OpenStudy (anonymous):

$5400  is invested, part of it at  11%  and part of it at  6%.  For a certain year, the total yield is  $ 474.00.   How much was invested at each rate? @Directrix

Directrix (directrix):

Let x be the amount invested at 11% Then, (5400 - x) is invested at 6% @dontre

Directrix (directrix):

.11*x + .06 * (5400 - x) = 474.00

OpenStudy (anonymous):

59400

Directrix (directrix):

.11*x + .06 * (5400 - x) = 474.00 .11*x + 324 -.06*x = 474,00

OpenStudy (anonymous):

confused again...I am trying to ge the hang of this

Directrix (directrix):

.05*x + 324 = 474 x = 3000 at 11% (5400 - x) is invested at 6% What is (5400 - x) @dontre

OpenStudy (anonymous):

0

Directrix (directrix):

No way. You are making up numbers.

Directrix (directrix):

What is (5400 - 3000) ?

Directrix (directrix):

3000 at 11% ? is invested at 6%

OpenStudy (anonymous):

3000 at 11% is 330

Directrix (directrix):

^^^^ is not the question.

Directrix (directrix):

This is the question:  How much was invested at each rate?

Directrix (directrix):

Answers: 3000 at 11% 2400 at 6%

OpenStudy (anonymous):

I just suck...been at this all day

OpenStudy (anonymous):

18000....1800